### v1

# New polarimetric study of the galactic X-ray burster GX 13$+$1

11institutetext: Department of Physics and Astronomy, FI-20014 University of Turku, Finland
11email: anna.a.bobrikova@utu.fi
22institutetext: INAF Istituto di Astrofisica e Planetologia Spaziali, Via del Fosso del Cavaliere 100, 00133 Roma, Italy 33institutetext: Dipartimento di Fisica, Università degli Studi di Roma “Tor Vergata”, Via della Ricerca Scientifica 1, 00133 Roma, Italy 44institutetext: Dipartimento di Fisica, Università degli Studi di Roma “La Sapienza”, Piazzale Aldo Moro 5, 00185 Roma, Italy
New polarimetric study of the galactic X-ray burster GX 13$+$1
Anna Bobrikova New polarimetric study of the galactic X-ray burster GX 13$+$1New polarimetric study of the galactic X-ray burster GX 13$+$1    Alessandro Di Marco New polarimetric study of the galactic X-ray burster GX 13$+$1New polarimetric study of the galactic X-ray burster GX 13$+$1    Fabio La Monaca New polarimetric study of the galactic X-ray burster GX 13$+$1New polarimetric study of the galactic X-ray burster GX 13$+$1New polarimetric study of the galactic X-ray burster GX 13$+$1New polarimetric study of the galactic X-ray burster GX 13$+$1New polarimetric study of the galactic X-ray burster GX 13$+$1New polarimetric study of the galactic X-ray burster GX 13$+$1    Juri Poutanen New polarimetric study of the galactic X-ray burster GX 13$+$1New polarimetric study of the galactic X-ray burster GX 13$+$1    Sofia V. Forsblom New polarimetric study of the galactic X-ray burster GX 13$+$1New polarimetric study of the galactic X-ray burster GX 13$+$1    Vladislav Loktev New polarimetric study of the galactic X-ray burster GX 13$+$1New polarimetric study of the galactic X-ray burster GX 13$+$1
(Received April 2, 2024; accepted XXX)
Weakly magnetized neutron stars (WMNS) are complicated sources with challenging phenomenology. For decades, they have been studied via spectrometry and timing. It has been established that the spectrum of WMNSs consists of several components traditionally associated with the accretion disk, the boundary or spreading layer, and the wind and their interactions with each other. Since 2022, WMNSs have been actively observed with the Imaging X-ray Polarimetry Explorer (IXPE). Polarimetric studies provided new information about the behavior and geometry of these sources. One of the most enigmatic sources of the class, galactic X-ray burster GX 13$+$1 was first observed with IXPE in October 2023. A strongly variable polarization at the level 2–5% was detected with the source showing a rotation of the polarization angle (PA) that hinted towards the misalignment within the system. The second observation was performed in February 2024 with a complementary observation by Swift/XRT. IXPE measured an overall polarization degree (PD) of 2.5% and the PA of 24°, and the Swift/XRT data helped us evaluate the galactic absorption and fit the continuum. Here we study the similarities and differences between the polarimetric properties of the source during the two observations. We confirm the expectation of the misalignment in the system and the assignment of the harder component to the boundary layer. We emphasize the importance of the wind in the system. We note the difference in the variation of polarimetric properties with energy and with time.
Key Words.:
accretion, accretion disks – polarization – stars: neutron – X-rays: binaries
1 Introduction
Although weakly magnetized neutron stars (WMNS) do not show pulsation or strong orbital variation, they present a complex phenomenology that requires further studies and observations to be explained. The broad class of WMNSs includes sources with a magnetic field of the order of $10^{7}-10^{9}$ G. These neutron stars accrete matter from the companion via the Roche lobe, and if the accretion rate is high enough to overpower the relatively weak magnetic field, the matter from the accretion disk falls directly onto the equator of the neutron star.
Non-pulsating weakly magnetized neutron stars are traditionally classified into two main groups: Z and atolls, on the basis of the patterns they show on the color-color diagram (CCD) or hardness-intensity diagram (HID). Although the classification is completely phenomenological, objects of these two classes have many other characteristic behavioral properties. For instance, atolls are known to be less bright than the Z-sources: typical luminosity of the atolls is in the range of $10^{36}-10^{37}$ erg s-1, while Z-sources are among the brightest objects in the X-ray sky with luminosities on the order of $10^{38}$ erg s-1. However, many features are known to be present in both classes: for example, quasi-periodic oscillations (QPOs) in the Hz and kHz ranges were observed from both Z and atolls sources (van der Klis, 1989, 2000). Sources of both classes are observed in both hard and soft states, and the main components of their spectrum are expected to be the same. The spectrum is generally well described by the two main components (see e.g. Revnivtsev et al., 2013). The soft component can be approximated by the multicolor black body emission of temperature $<1$ keV associated with the accretion disk. The harder component most probably comes from Comptonization in a relatively cool plasma (2–3 keV), either a boundary layer between the accretion disk and the surface of the neutron star (Shakura & Sunyaev, 1988), or the spreading layer at the neutron star surface (Inogamov & Sunyaev, 1999; Suleimanov & Poutanen, 2006). Additional features known to the spectrum are the broad Fe emission line associated with the reflection of the spreading layer emission from the accretion disk and sharp absorption lines coming from the absorption of the emission in the wind above the disk.
However, the geometry of the WMNS cannot be fully understood from spectroscopic and time-resolved observations only. Information obtained from variability, and specifically from the QPOs, revealed that the strongest variability is associated with a harder spectral component (Gilfanov et al., 2003; Revnivtsev & Gilfanov, 2006; Revnivtsev et al., 2013) originating close to the neutron star surface. Polarimetry brought a new perspective to the problem. For instance, it became possible to distinguish between the boundary and spreading layer emission based on the polarimetric properties of the emission, while spectroscopically they are identical. The boundary layer is expected to have a polarization similar to that of the accretion disk, with the polarization vector lying in the plane of the disk (Chandrasekhar, 1960; Loktev et al., 2022), while the polarization of the spreading layer is expected to be perpendicular to the accretion disk (Farinelli et al., 2024). For example, the polarization angle (PA) of the bright Z-source Cyg X-2 (Farinelli et al., 2023) turned out to be consistent with the position angle of the radio jet (which is likely perpendicular to the accretion disk), implying the origin of the polarization in the spreading layer or from scattering in the accretion disk wind (Tomaru et al. 2024; Nitindala et al., in prep.). In La Monaca et al. (2024), on the other hand, a significant difference between the jet position angle and the PA for Sco X-1 has been observed which implies a more complicated geometry of the source. In the peculiar case of Cir X-1 (Rankin et al., 2024), the variability of the PA with time and hardness supported the idea of a misalignment between the angular momentum of the neutron star and the binary orbital axis. In Z-sources XTE J1701--462 (Cocchi et al., 2023) and GX 5--1 (Fabiani et al., 2023), a significant variation of the polarization between observations in the hard and soft states suggested the change in geometry between states. In atoll sources GX 9+9 (Ursini et al., 2023), 4U 1820--303 (Di Marco et al., 2023a), and 4U 1624--49 (Saade et al., 2024), the strong dependency of the PD on energy supported the separation of the spectrum into two components with the different PAs. These discoveries supported the study of the WMNSs and established the capabilities of polarimetric observations of these sources.
Galactic X-ray burster GX 13$+$1 was first observed by IXPE on 2023 October 17–19 (Bobrikova et al., 2024, hereafter B24). This source is known for the strong wind above the large disk being responsible for a number of absorption lines (Díaz Trigo et al., 2012), the brightness of $0.5L_{\rm Edd}$ (D’Aí et al., 2014), and an orbital period of 24.5 days, which is very long for a low-mass X-ray binary (Corbet et al., 2010). GX 13$+$1 presents a classification challenge: it is still debatable whether the source is a Z (see e.g. Saavedra et al., 2023) or an atoll (see e.g. Schnerr et al., 2003).
During the first IXPE observation, GX 13$+$1 presented an enigmatic behavior: the polarimetric properties varied significantly during the observation, while the spectral properties remained almost unchanged. Moreover, GX 13$+$1 shows a peculiar continuous rotation of the PA by 70° in the two days of observations, together with a change in the pattern of dependence of PD on energy. These two peculiar behaviors left some space for further investigation. Here, we aim to answer some of the open questions and bring a new perspective to the study of GX 13$+$1.
The remainder of the paper is structured as follows. We introduce the observations performed with IXPE and Swift/XRT observatories in Sect. 2. In Sect. 3 we present the data analysis. We provide possible interpretations of the results in Sect. 4 and a summary in Sect. 5.
2 Observations
2.1 IXPE
IXPE is a joint mission of NASA and the Italian Space Agency (ASI) that was launched on 9 December 2021. The observatory comprises three grazing incidence mirror assembly modules, each having a detector unit (DU), hosting an X-ray detector sensitive to polarization, the Gas Pixel Detector (GPD), in the focal plane (Costa et al., 2001; Baldini et al., 2021). A description of the mission and of the instrument on board is given in Soffitta et al. (2021), and Weisskopf et al. (2022).
IXPE observed GX 13$+$1 two times, the first one performed in October 2023, reported in B24, and the present standing since 25 February 2024 at 13:34 UTC until 27 February at 12:59 UTC for a total exposure time of about 90 ks in each DU (see Table 1 and the light curve in Fig. 1). Data have been processed for model-independent polarimetric analysis using the ixpeobssim package version 30.2.1 (Baldini et al., 2022). Spectral and spectro-polarimetric analysis was performed using HEASoft version 6.33 and the standard FTOOLS (Nasa High Energy Astrophysics Science Archive Research Center (2014), Heasarc) with the IXPE CALDB released on 28 February 2024. Source photons were selected in a circular region with a radius of 100″  centered at the source position. As recommended by Di Marco et al. (2023b), given the brightness of the source, the background was not subtracted. The unweighted analysis is performed with ixpeobssim, while the weighted analysis is adopted for the spectropolarimetric analysis as suggested by Di Marco et al. (2022a).
Table 1: List of IXPE and Swift/XRT observations: Observation ID, start, stop, and the exposure times are reported.
Observatory ObsID Start – Stop Exposure Time (s)
Swift/XRT-XRT 00036688042 2024-02-26T05:05–2024-02-26T05:27 1366
00036688043 2024-02-25T13:142024-02-25T17:59 892
00036688044 2024-02-26T13:07–2024-02-26T14:52 1249
00036688045 2024-02-25T22:53–2024-02-26T00:35 1395
00036688046 2024-02-26T17:56–2024-02-26T19:30 1311
00036688047 2024-02-26T08:22–2024-02-26T10:12 1429
00036688048 2024-02-25T19:28–2024-02-25T21:32 1321
00036688049 2024-02-26T22:24–2024-02-26T23:59 771
00036688050 2024-02-27T03:22–2024-02-27T05:14 507
00036688051 2024-02-27T11:05–2024-02-27T11:29 1443
IXPE 03001101 2024-02-25T13:34–2024-02-27T12:59 90670
90819
90799
Figure 1: The IXPE light curve obtained combining the 3 DUs (top panel) and the corresponding hardness ratio as a function of time (bottom panel). The data are binned in 200 s. The shaded regions mark the Swift/XRT observation periods. Dashed vertical lines illustrate the separation of the observation into five equal time bins of 9.5 h.
Figure 2: Hardness-Intensity diagram obtained from IXPE data. The data are binned in 200 s.
2.2 Swift/XRT
The Neil Gehrels Swift Observatory (Gehrels et al., 2004) carrying three instruments on board, enables the most detailed observations of gamma-ray bursts to date; the X-ray Telescope (XRT) is one of the instruments on board based on a sensitive, flexible, autonomous X-ray CCD imaging spectrometer. Swift/XRT coordinated observations with IXPE for GX 13$+$1 were performed; they are used to monitor the status of the source and obtain spectral information during IXPE observation. Given the source brightness, the Swift/XRT observations were performed in Windowed Timing mode (WT). Ten Swift/XRT pointings covered the IXPE observation, as reported in Table 1 and Fig. 1. Swift/XRT data were extracted using HEASoft v6.33 and standard FTOOLS Nasa High Energy Astrophysics Science Archive Research Center (2014) (Heasarc). Source and background extractions were performed using Swift/XRT imaging capabilities in an annulus with an inner radius of about 5″ and an outer radius of about 60″ to avoid possible pile-up effects (Romano et al., 2006). The Swift/XRT data have been fitted in the energy band 0.7–8 keV to have sufficient statistics in the spectral fits and grouped to have at least 50 counts bin-1. In the analysis, the response matrices released in the HEASARC CALDB on 2023 July 25 have been applied.
3 Data analysis
IXPE observed GX 13$+$1 for $\sim$90 ks. During the observation, as shown in Fig. 1 (top panel), the source rate was rather stable, with a slight increase at the end of the observation. The source hardness in the IXPE data appears to have no strong variations, remaining almost constant throughout the observation; this is in contrast to the previous IXPE observation, reported in B24, where a dip corresponding to a slight hardening was present. In Fig. 2, the hardness intensity diagram obtained from the IXPE data is reported and shows that hardness did not change by more than 10%.
Swift/XRT observations, covering with short snapshots the IXPE observation, allow us to monitor the source spectral state. GX 13$+$1 studies in the past showed a spectrum that can be well described, as typical in the atoll and Z sources, by a soft diskbb component describing disk emission and Comptonization emission from the spreading/boundary layer that can be described by a bbodyrad component (Díaz Trigo et al., 2012). Better modeling of the hard Comptonization component, e.g., applying compTT or compTB, is not possible due to the limited IXPE and Swift/XRT energy bands. In the literature, XMM–Newton spectra of GX 13$+$1 are reported showing wind and reflection features (Díaz Trigo et al., 2012), but neither IXPE nor Swift/XRT allow us to study these features, so in the following the simple spectral model tbabs*(diskbb+bbodyrad) will be applied.
3.1 Spectroscopy
Given the better spectral capabilities, we used Swift/XRT data as a driver for spectral modeling and to monitor spectral variations. The analysis is carried out in the energy band 0.7–8 keV, best-fit values are reported for each observation in Table 2 with uncertainties at the 68% confidence level (CL), and ObservationIDs are ordered by the observation date to show the time evolution of the spectral model.
Table 2: Best-fit parameters for the spectral model TBabs*(diskbb+bbodyrad) as obtained by the different Swift/XRT spectra.
[…]
$N_{\text{H}}$ (1022 cm2) $kT_{\text{in}}$ (keV) norm $kT_{\text{bb}}$ (keV) norm (10-9 erg cm-2 s-1)
00036688043 $4.54^{+0.18}_{-0.14}$ $1.1^{+0.6}_{-0.2}$ $250^{+260}_{-160}$ $1.5^{+0.6}_{-0.1}$ $100^{+60}_{-90}$ 0.93 6.0 0.35
00036688048 $4.52\pm 0.10$ [1.0] $290\pm 19$ $1.52\pm 0.04$ $120\pm 12$ 0.91 5.9 0.37
00036688045 $5.71\pm 0.19$ $0.67\pm 0.05$ $1100\pm 400$ $1.57\pm 0.04$ $100\pm 10$ 1.13 4.1 0.45
00036688042 $5.0\pm 0.2$ $0.70^{+0.10}_{-0.07}$ $1000^{+700}_{-400}$ $1.41^{+0.06}_{-0.04}$ $150\pm 30$ 1.06 5.9 0.38
00036688047 $4.62_{-0.18}^{+0.14}$ $0.88_{-0.10}^{+0.38}$ $500\pm 300$ $1.38^{+0.18}_{-0.05}$ $200_{-100}^{+40}$ 1.12 6.0 0.36
00036688044 $4.67_{-0.15}^{+0.17}$ $0.98_{-0.15}^{+0.39}$ $400^{+300}_{-200}$ $1.46^{+0.31}_{-0.09}$ $130_{-100}^{+60}$ 1.04 5.9 0.33
00036688046 $4.9_{-0.2}^{+0.3}$ $0.70^{+0.11}_{-0.08}$ $1100_{-600}^{+900}$ $1.29_{-0.04}^{+0.06}$ $230^{+40}_{-50}$ 1.10 5.0 0.33
00036688049 $4.3\pm 0.3$ $1.0_{-0.2}^{+0.6}$ $320_{-260}^{+630}$ $1.7_{-0.1}^{+1.3}$ $80_{-75}^{+65}$ 1.16 5.4 0.34
00036688050 $5.7\pm 0.4$ $0.66^{+0.12}_{-0.09}$ $1800_{-1000}^{+2000}$ $1.46_{-0.08}^{+0.11}$ $130\pm 40$ 0.92 4.7 0.36
00036688051 $4.70_{-0.09}^{+0.12}$ $1.13^{+0.19}_{-0.16}$ $310_{-120}^{+200}$ $1.77_{-0.17}^{+0.28}$ $80_{-40}^{+50}$ 1.16 7.9 0.35
 [^1] The fits are performed in the energy band 0.7–8.0 keV. Errors correspond to the 68% CL. The reported flux has a typical uncertainty of the order of $0.3\times 10^{-9}$. The hardness is defined as in Eq. (1).
In the spectral analysis, the Swift/XRT data are fitted with the model: tbabs*(diskbb+bbodyrad). In the different pointings, it is possible to observe that the absorption is in average $\sim 4.7\times 10^{22}$ cm-2, while the disk temperature ranges between 0.7–1.1 keV, about the bbodyrad component – which in our case is used to describe a Comptonization emission from the boundary layer – the temperature ranges in the $\sim 1.3-1.8$ keV. The flux is in the range $\sim$(4–6)×$\times 10^{-9}$ erg cm-2 s-1 except in the last pointing where it reaches the highest value of $\sim 8\times 10^{-9}$ erg cm-2 s-1. The hardness, reported in Table 2 is defined as
$\text{Hardness}=\frac{\text{Flux}_{0.7-3\leavevmode\nobreak\ \text{keV}}-\text{Flux}_{3-8\leavevmode\nobreak\ \text{keV}}}{\text{Flux}_{0.7-8\leavevmode\nobreak\ \text{keV}}}$ (1)
and the spectral models report an almost constant hardness value except for the case of ObservationID 00036688045 where the flux is at a minimum and there is a slight hardening, but this Swift/XRT observation (third snapshot) is performed mainly during an IXPE occultation period. Thus, the IXPE data are not affected by this slight hardening, while the other snapshots confirm the IXPE result of an almost constant hardness ratio during observation. In the following analysis, for the spectro-polarimetric study, this ObservationID will be excluded, while the other ones will be used in the joint fit with the IXPE data.
Taking into account the whole IXPE observation, and the Swift/XRT spectra, except the ObservationID 00036688045, we obtain a good joint spectrum model. The best-fit parameters are reported in Table 3 and Fig. 3. To take into account the calibration uncertainties in the IXPE data, (see, e.g., Di Marco et al., 2022b), we left the gain slope and offset of the IXPE-DUs free to vary.
Table 3: Best-fit parameters of the model const*tbabs*(diskbb+bbodyrad) for the joint Swift/XRT and IXPE spectra.
Model Parameter Value
tbabs $N_{\text{H}}$ (1022 cm2) $4.45_{-0.05}^{+1.0}$
diskbb $kT_{\text{in}}$ (keV) $0.88_{-0.07}^{+0.10}$
norm $390^{+140}_{-130}$
bbodyrad $kT_{\text{bb}}$ (keV) $1.33^{+0.13}_{-0.14}$
norm $228_{-8}^{+20}$
const Swift/XRT/XRT [1]
IXPE-DU1 $0.722\pm 0.003$
IXPE-DU2 $0.686\pm 0.003$
IXPE-DU3 $0.658\pm 0.003$
$\chi^{2}$/d.o.f. 991/956=1.04
Flux ($10^{-9}$erg cm-2 s-1)
2–8 keV 6.0
diskbb 1.3
bbodyrad 4.7
 [^2] Errors correspond to the 68% CL.
Figure 3: Joint spectral fit Swift/XRT and IXPE in the 1–8 keV energy range. In the top panel the spectra of each telescope are reported in $EF_{E}$ units, in the bottom panel the residuals are shown.
The joint fit confirms the result expected for the average model from several Swift/XRT snapshots. In the following for the spectro-polarimetric analysis, the spectral model will be frozen to the present one. The IXPE gain correction is at level of 0.95 for the slope in all the three DUs and the offset in the range 70–100 eV.
3.2 Polarimetry
Spectroscopy revealed that neither the light curve nor the hardness have strong dips or features that would suggest separating the observation into several parts to be studied independently. To study the time variability of the polarimetric properties, we followed the same approach as in B24 and split this new IXPE observation into five equal time bins spanning 9.5 h. The results are reported in Fig. 4.
Figure 4: Time dependence of the normalized Stokes parameters $q$ and $u$ (top), and of the PD and PA (bottom) obtained by the pcube algorithm using the data separated into 9.5 h time bins. The confidence regions are reported at $68\%$.
Figure 5: Energy resolved normalized Stokes parameters $q$ and $u$ (top), and PD and PA (bottom) obtained by the pcube algorithm using the reported energy bins for the time-averaged data. The confidence regions are reported at $68\%$ CL. The vectors correspond to the best-fit polarization parameters of the diskbb and bbodyrad components, respectively, as obtained in the spectro-polarimetric analysis and reported in Table 4.
[…]
3.3 Spectro-polarimetry
[…]
Table 4: IXPE spectro-polarimetric fit of the Stokes parameters $I$, $Q$, and $U$ using the spectral model from Table 3.
Model PD/$A_{1}$ (%) $A_{\rm slope}$ (keV-1) PA/$\psi_{1}$ (deg) $\psi_{\rm slope}$ (deg,keV-1) $\chi^{2}$/d.o.f.
tbabs*(diskbb+bbodyrad)*polconst $2.4\pm 0.3$ $28\pm 3$ 243/250=0.97
tbabs*(diskbb+bbodyrad)*pollin $3.8\pm 0.8$ $-0.5\pm$0.3 $28\pm 3$ $-8\pm$4 234/248=0.94
tbabs*(diskbb*polconst $6.5\pm 1.6$ $41\pm 7$ 235/248=0.95
+bbodyrad*polconst) $1.6\pm 0.7$ $3\pm 13$
 [^3] Errors correspond to 68% CL.
Figure 6: Spectral energy distribution of GX 13$+$1 in $EF_{E}$ representation as observed by IXPE. The left, middle, and right panels are for the Stokes parameters $I$, $Q$, and $U$, respectively. The fit is performed in the 2–8 keV energy band using the three IXPE detectors and applying the model tbabs*(diskbb*polconst+bbodyrad*polconst). The total model is shown with the solid black line, while the diskbb and bbodyrad with the dashed and dotted lines, respectively. The lower sub-panels show the residuals between the data and the best fit. In the joint spectral fit, we apply a rebinning to have at least 50 counts per bin as in Swift/XRT, while for the spectro-polarimetric analysis, a constant grouping is applied to have energy bins of 200 eV.
4 Discussion
Although interesting as a single observation, the results reported in this article provide the most information when compared to the results of the October observations reported in B24. Overall, in October IXPE measured a PD of $\sim$1.4and a PA of $\sim-2\degr$, significantly different from the current results. However, the reason for the depolarization in the first observation was fast variability in the PA. In B24, the observation was separated into the pre-dip, dip, and post-dip states. The results for the dip were poorly constrained, but the other two are presented in Fig. 7 and compared to the results of the current observation. We note that the current observation is similar in PA to the post-dip state of B24. The lower PD in the entire current observation most probably comes from averaging the entire observation in time and energy, as for the previous one. In Fig. 4, the Bin 3, for instance, shows a polarization within errors similar to the one measured after the dip in B24.
[…]
Both HID and the shape of the spectrum confirm that GX 13$+$1 was in a soft state throughout the observation. We do not see any significant change of state in the IXPE and Swift/XRT energy bands, although, in the last part of the observation, the flux rises, which is a sign of the source going into an even softer state. If we look at the HID of the source in the October observation (B24, Figure 2), we see that the source is almost in the same state, except for the dipping period. In this new observation, since simultaneous Swift/XRT observations offer a better spectral capability than IXPE, the spectral modeling and decomposition into two main components are better determined, allowing for a better spectro-polarimetric analysis.
The light curve does not show any peculiar variability: unlike the one from the October observation, the current one has no strong dips or rapid changes, and neither does the hardness ratio. Despite this constant hardness, we attempted to study the possible variation in 9.5 h equal time bins similarly to B24. This analysis reports a hint for a variation in PD in Fig. 4, in particular is interesting that in Bin 5, corresponding to a softening of GX 13$+$1 there is a lower polarization and an indication for a possible rotation with respect to the other 4 time bins. The clear rotation of PA observed during the first observation is not present this time, but the PA appears to have remained aligned with the post-dip polarization measured in B24. If we compare the positions of the points on the $(q,u)$ plane with those obtained from the October 2023 observation (Figure 7 in B24), we see a different pattern, but the points are still in the I or IV quadrant. The biggest difference is in order: in October, we saw the points following the straight line path, but here the variability shows a different pattern.
[…]
From the aligned systems, we expect the polarization vectors of the emission coming from the disk and the Comptonized component to be either nearly parallel or perpendicular to each other. Hence, we expect some misalignment in the system that could support this difference in the PA. Another thing to notice is a very high PD of the disk component: as was mentioned in Loktev et al. (2022), we do not expect a polarization higher than 4% in the IXPE energy range even for the highest inclination. This value is compatible with the spectro-polarimetric results at 90% CL. However, knowing that the source is likely to have an inclination lower than 80°(Díaz Trigo et al., 2012; Tomaru et al., 2020), we need to add the scattering of the disk emission in the wind to reach the PD of 6.5%. We assume that the Comptonized component comes from the boundary layer, as it is expected to align in PA with the disk emission, and the difference in PAs is less than 45°.
In search of a source of misalignment, we looked at the previous observations of IXPE. It is possible (although unexpected) that, similar to the Cir X-1 (Rankin et al., 2024), the rotation axis of the neutron star is slightly misaligned from the rotation axis of the binary system. The boundary layer would then be dragged out of the disk plane by the motion of the star. Another option would be the peculiar geometry of the wind in the system.
Finally, we want to address the significant difference between the two observations of GX 13$+$1. The object was in the same spectral state but half an orbit away from the previous observation, when the source clearly showed a rotation of the PA with time. However, in February 2024, there was no clear monotonic change in the PA, while we still see variations in the PA. One possibility is that the mechanism responsible for the dip in the light curve during the first observation was also responsible for the rotation of the PA. However, this mechanism is unknown. What we can address is that from this new observation, it appears that the PA remained aligned with the one observed in the end of the previous observation. Then, at the end of the present IXPE pointing, when the flux grows and the source goes into a softer state, the PD decreases and rotates to a PA similar to the one we observed in October before the dip. This could suggest, similarly to what we observed in Cir X-1, a variation of the PA with the state of the source. This could mean a variation of the geometry of the hot region in different states; this can also be confirmed by the indication we obtained in the October 2023 observation when the energy dependence of the polarization changed before and after the dip going from an energy-dependent behavior, as observed e.g. for 4U 1820--30 (Di Marco et al., 2023a), into a constant polarization with energy, as observed, e.g., for Sco X-1 in La Monaca et al. (2024). Present data do not allow for a clear statement on this, but we hope in the future we could obtain further observations to improve the significance for a detailed analysis of the polarization along the states of atoll and Z sources.
A further point is the contribution of the wind, as we are unaware of the wind behavior during both observations, we can also assume that it could be responsible for the differences in polarimetry, yet we have no way to support this claim.
5 Summary
We analyzed the observation of GX 13$+$1 carried out simultaneously by IXPE and Swift/XRT. In this article, we report the highly significant detection of polarization from this source in the soft state. Spectroscopic analysis showed that the spectrum is close to that in Díaz Trigo et al. (2012), and the hardness of the emission confirmed the lower left banana state, as well as the further softening of the spectrum at the end of this new observation. We used a diskbb model for the softer disk emission, and the bbodyrad model to approximate the harder component associated with the spreading or boundary layer. The polarimetric analysis provided the overall PD of the source during the observation at 2.5% at ¿$5\sigma$ CL, and a PA at 24°. We also saw a marginally significant decrease in PD and PA with energy. The same trend was confirmed in the spectro-polarimetric analysis, which showed that the softer component is stronger polarized than the harder component with the PA differing by $\sim$40°. The results of the spectro-polarimetric analysis using xspec performed under the assumption of one overall polarization model agreed with the pcube polarimetric results.
We studied the differences and similarities between the first (B24) and the second observations of GX 13$+$1 by IXPE. Our main conclusion is that we saw a significantly higher overall polarization in the second observation, even though the source was in the same state during both observations. The depolarization comes from the variability of polarimetric properties, which was larger in the first observation. We discussed possible reasons for the different polarimetric behavior in the two observations.
Acknowledgements.
This research used data products provided by the IXPE Team (MSFC, SSDC, INAF, and INFN) and distributed with additional software tools by the High-Energy Astrophysics Science Archive Research Center (HEASARC), at NASA Goddard Space Flight Center (GSFC). The Imaging X-ray Polarimetry Explorer (IXPE) is a joint US and Italian mission. The Italian contribution is supported by the Italian Space Agency (Agenzia Spaziale Italiana, ASI) through contract ASI-OHBI-2022-13-I.0, agreements ASI-INAF-2022-19-HH.0 and ASI-INFN-2017.13-H0, and its Space Science Data Center (SSDC) with agreements ASI-INAF-2022-14-HH.0 and ASI-INFN 2021-43-HH.0, and by the Istituto Nazionale di Astrofisica (INAF) and the Istituto Nazionale di Fisica Nucleare (INFN) in Italy. We thank the Swift Project Scientists for approving our DDT request to observe GX 13$+$1. This research has been supported by the Finnish Cultural Foundation grant 00240328 (AB) and the Academy of Finland grant 333112 (AB, JP, SVF, VL). ADM, FLM are partially supported by MAECI with grant CN24GR08 “GRBAXP: Guangxi-Rome Bilateral Agreement for X-ray Polarimetry in Astrophysics”.

### v2

@@
 (Received April 2, 2024; accepted XXX)
-Weakly magnetized neutron stars (WMNS) are complicated sources with challenging phenomenology.
-For decades, they have been studied via spectrometry and timing.
-It has been established that the spectrum of WMNSs consists of several components traditionally associated with the accretion disk, the boundary or spreading layer, and the wind and their interactions with each other.
-Since 2022, WMNSs have been actively observed with the Imaging X-ray Polarimetry Explorer (IXPE).
-Polarimetric studies provided new information about the behavior and geometry of these sources.
-One of the most enigmatic sources of the class, galactic X-ray burster GX 13$+$1 was first observed with IXPE in October 2023.
-A strongly variable polarization at the level 2–5% was detected with the source showing a rotation of the polarization angle (PA) that hinted towards the misalignment within the system.
-The second observation was performed in February 2024 with a complementary observation by Swift/XRT.
-IXPE measured an overall polarization degree (PD) of 2.5% and the PA of 24°, and the Swift/XRT data helped us evaluate the galactic absorption and fit the continuum.
-Here we study the similarities and differences between the polarimetric properties of the source during the two observations.
-We confirm the expectation of the misalignment in the system and the assignment of the harder component to the boundary layer.
-We emphasize the importance of the wind in the system.
-We note the difference in the variation of polarimetric properties with energy and with time.
+Weakly magnetized neutron stars (WMNSs) are complex astrophysical objects with challenging phenomenology.
+For decades, they have been studied via spectrometry and timing analyses.
+It is well established that the spectrum of WMNSs consists of several components traditionally associated with the accretion disk, the boundary or spreading layer, and the wind, along with their interactions.
+Since 2022, WMNSs have been actively observed using the Imaging X-ray Polarimetry Explorer (IXPE).
+Polarimetric studies have provided new information about the behavior and geometry of these sources.
+One of the most enigmatic sources in this category, the galactic X-ray burster GX 13$+$1, was first observed with IXPE in October 2023.
+A highly variable polarization at levels of 2–5% was detected, with the source showing a rotation of the polarization angle (PA), suggestive of misalignment within the system.
+A second observation was performed in February 2024, complemented by observations from Swift/XRT.
+IXPE measured an overall polarization degree (PD) of 2.5% and a PA of 24°, while Swift/XRT data helped us evaluate the galactic absorption and fit the continuum.
+Here, we study the similarities and differences in the polarimetric properties of the source during the two observations.
+Our findings confirm the expected misalignment in the system and the assignment of the harder component to the boundary layer.
+We also emphasize the significance of the wind in the system.
+Additionally, we observe notable differences in the variation of polarimetric properties with energy and over time.
 Key Words.:
@@
 1 Introduction
-Although weakly magnetized neutron stars (WMNS) do not show pulsation or strong orbital variation, they present a complex phenomenology that requires further studies and observations to be explained.
-The broad class of WMNSs includes sources with a magnetic field of the order of $10^{7}-10^{9}$ G.
-These neutron stars accrete matter from the companion via the Roche lobe, and if the accretion rate is high enough to overpower the relatively weak magnetic field, the matter from the accretion disk falls directly onto the equator of the neutron star.
-Non-pulsating weakly magnetized neutron stars are traditionally classified into two main groups: Z and atolls, on the basis of the patterns they show on the color-color diagram (CCD) or hardness-intensity diagram (HID).
+Although weakly magnetized neutron stars (WMNSs) do not show pulsation or strong orbital variation, they present a complex phenomenology that requires further studies and observations to be explained.
+The broad class of WMNSs includes sources with a magnetic field on the order of $10^{7}-10^{9}$ G.
+These neutron stars accrete matter from their companions via the Roche lobe.
+If the accretion rate is high enough to overpower the relatively weak magnetic field, the matter from the accretion disk falls directly onto the equator of the neutron star.
+Non-pulsating WMNSs are traditionally classified into two main groups, Z and atolls, based on the patterns they show on the color-color diagram (CCD) or hardness-intensity diagram (HID).
 Although the classification is completely phenomenological, objects of these two classes have many other characteristic behavioral properties.
-For instance, atolls are known to be less bright than the Z-sources: typical luminosity of the atolls is in the range of $10^{36}-10^{37}$ erg s-1, while Z-sources are among the brightest objects in the X-ray sky with luminosities on the order of $10^{38}$ erg s-1.
-However, many features are known to be present in both classes: for example, quasi-periodic oscillations (QPOs) in the Hz and kHz ranges were observed from both Z and atolls sources (van der Klis, 1989, 2000).
+For instance, atolls are known to be less bright than the Z sources: the typical luminosity of the atolls is in the range of $10^{36}-10^{37}$ erg s-1, while Z sources are among the brightest objects in the X-ray sky, with luminosities on the order of $10^{38}$ erg s-1.
+However, many features are known to be present in both classes.
+For example, quasi-periodic oscillations (QPOs) in the hertz and kilohertz ranges were observed from both Z and atoll sources (van der Klis, 1989, 2000).
 Sources of both classes are observed in both hard and soft states, and the main components of their spectrum are expected to be the same.
-The spectrum is generally well described by the two main components (see e.g.
-Revnivtsev et al., 2013).
+The spectrum is generally well described by the two main components (see, e.g., Revnivtsev et al., 2013).
 The soft component can be approximated by the multicolor blackbody emission of temperature $<1$ keV associated with the accretion disk.
-The harder component most probably comes from Comptonization in a relatively cool plasma (2–3 keV), either a boundary layer between the accretion disk and the surface of the neutron star (Shakura & Sunyaev, 1988), or the spreading layer at the neutron star surface (Inogamov & Sunyaev, 1999; Suleimanov & Poutanen, 2006).
-Additional features known to the spectrum are the broad Fe emission line associated with the reflection of the spreading layer emission from the accretion disk and sharp absorption lines coming from the absorption of the emission in the wind above the disk.
-However, the geometry of the WMNS cannot be fully understood from spectroscopic and time-resolved observations only.
+The harder component is probably caused by Comptonization in a relatively cool plasma (2–3 keV), either a boundary layer between the accretion disk and the surface of the neutron star (Shakura & Sunyaev, 1988) or the spreading layer at the neutron star surface (Inogamov & Sunyaev, 1999; Suleimanov & Poutanen, 2006).
+Additional features known to the spectrum are the broad Fe emission line associated with the reflection of the spreading layer emission from the accretion disk and sharp absorption lines caused by the absorption of the emission in the wind above the disk.
+However, the geometry of the WMNSs cannot be fully understood from spectroscopic and time-resolved observations alone.
 Information obtained from variability, and specifically from the QPOs, revealed that the strongest variability is associated with a harder spectral component (Gilfanov et al., 2003; Revnivtsev & Gilfanov, 2006; Revnivtsev et al., 2013) originating close to the neutron star surface.
 Polarimetry brought a new perspective to the problem.
-For instance, it became possible to distinguish between the boundary and spreading layer emission based on the polarimetric properties of the emission, while spectroscopically they are identical.
+For instance, it became possible to distinguish between the boundary and spreading layer emission based on the polarimetric properties of the emission, while spectroscopically, they are identical.
 The boundary layer is expected to have a polarization similar to that of the accretion disk, with the polarization vector lying in the plane of the disk (Chandrasekhar, 1960; Loktev et al., 2022), while the polarization of the spreading layer is expected to be perpendicular to the accretion disk (Farinelli et al., 2024).
-For example, the polarization angle (PA) of the bright Z-source Cyg X-2 (Farinelli et al., 2023) turned out to be consistent with the position angle of the radio jet (which is likely perpendicular to the accretion disk), implying the origin of the polarization in the spreading layer or from scattering in the accretion disk wind (Tomaru et al.
+For example, the polarization angle (PA) of the bright Z-source Cyg X-2 (Farinelli et al., 2023) turned out to be consistent with the position angle of the radio jet (which is likely perpendicular to the accretion disk), implying the polarization originated in the spreading layer or from scattering in the accretion disk wind (Tomaru et al.
 2024; Nitindala et al., in prep.).
 In La Monaca et al.
-(2024), on the other hand, a significant difference between the jet position angle and the PA for Sco X-1 has been observed which implies a more complicated geometry of the source.
-In the peculiar case of Cir X-1 (Rankin et al., 2024), the variability of the PA with time and hardness supported the idea of a misalignment between the angular momentum of the neutron star and the binary orbital axis.
-In Z-sources XTE J1701--462 (Cocchi et al., 2023) and GX 5--1 (Fabiani et al., 2023), a significant variation of the polarization between observations in the hard and soft states suggested the change in geometry between states.
-In atoll sources GX 9+9 (Ursini et al., 2023), 4U 1820--303 (Di Marco et al., 2023a), and 4U 1624--49 (Saade et al., 2024), the strong dependency of the PD on energy supported the separation of the spectrum into two components with the different PAs.
-These discoveries supported the study of the WMNSs and established the capabilities of polarimetric observations of these sources.
-Galactic X-ray burster GX 13$+$1 was first observed by IXPE on 2023 October 17–19 (Bobrikova et al., 2024, hereafter B24).
-This source is known for the strong wind above the large disk being responsible for a number of absorption lines (Díaz Trigo et al., 2012), the brightness of $0.5L_{\rm Edd}$ (D’Aí et al., 2014), and an orbital period of 24.5 days, which is very long for a low-mass X-ray binary (Corbet et al., 2010).
-GX 13$+$1 presents a classification challenge: it is still debatable whether the source is a Z (see e.g.
-Saavedra et al., 2023) or an atoll (see e.g.
-Schnerr et al., 2003).
+(2024), on the other hand, a significant difference between the jet position angle and the PA for Sco X-1 has been observed, which implies a more complicated geometry of the source.
+In the peculiar case of Cir X-1 (Rankin et al., 2024), the variability of the PA with time and hardness supports the idea of a misalignment between the angular momentum of the neutron star and the binary orbital axis.
+In Z-sources XTE J1701--462 (Cocchi et al., 2023) and GX 5--1 (Fabiani et al., 2024), a significant variation in the polarization between observations in the hard and soft states suggested a change in geometry between states.
+In atoll sources GX 9+9 (Ursini et al., 2023), 4U 1820--303 (Di Marco et al., 2023), and 4U 1624--49 (Saade et al., 2024), the strong dependence of the polarization degree (PD) on energy supports the separation of the spectrum into two components with the different PAs.
+These discoveries demonstrate the usefulness of using polarimetric observations to study these sources.
+The galactic X-ray burster GX 13$+$1 was first observed by Imaging X-ray Polarimetry Explorer (IXPE) on 2023 October 17–19 (Bobrikova et al., 2024, hereafter B24).
+This source is known for the strong wind above the large disk being responsible for several absorption lines (Díaz Trigo et al., 2012), a brightness of $0.5L_{\rm Edd}$ (D’Aí et al., 2014), and an orbital period of 24.5 d, which is very long for a low-mass X-ray binary (Corbet et al., 2010).
+GX 13$+$1 presents a classification challenge: it is still debatable whether it is a Z (see, e.g., Saavedra et al., 2023) or an atoll source (see, e.g., Schnerr et al., 2003).
 During the first IXPE observation, GX 13$+$1 presented an enigmatic behavior: the polarimetric properties varied significantly during the observation, while the spectral properties remained almost unchanged.
@@
 Here, we aim to answer some of the open questions and bring a new perspective to the study of GX 13$+$1.
-The remainder of the paper is structured as follows.
-We introduce the observations performed with IXPE and Swift/XRT observatories in Sect.
+The remainder of the paper is structured as follows: we introduce the observations performed with the IXPE and Swift/XRT observatories in Sect.
 2.
 In Sect.
-3 we present the data analysis.
+3, we present our data analysis.
 We provide possible interpretations of the results in Sect.
@@
 2.1 IXPE
-IXPE is a joint mission of NASA and the Italian Space Agency (ASI) that was launched on 9 December 2021.
-The observatory comprises three grazing incidence mirror assembly modules, each having a detector unit (DU), hosting an X-ray detector sensitive to polarization, the Gas Pixel Detector (GPD), in the focal plane (Costa et al., 2001; Baldini et al., 2021).
+IXPE is a joint mission of the National Aeronautics and Space Administration (NASA) and the Italian Space Agency (Agenzia Spaziale Italiana, ASI) that was launched on 2021 December 9.
+The observatory comprises three grazing-incidence mirror assembly modules, each of which has a polarization-sensitive X-ray detector unit (DU) in the focal plane (Costa et al., 2001; Baldini et al., 2021).
 A description of the mission and of the instrument on board is given in Soffitta et al.
-(2021), and Weisskopf et al.
+(2021) and Weisskopf et al.
 (2022).
-IXPE observed GX 13$+$1 two times, the first one performed in October 2023, reported in B24, and the present standing since 25 February 2024 at 13:34 UTC until 27 February at 12:59 UTC for a total exposure time of about 90 ks in each DU (see Table 1 and the light curve in Fig.
+IXPE observed GX 13$+$1 twice: the first observation was conducted in October 2023, as reported in B24, and the second took place from 2024 February 25, at 13:34 UTC to 2024 February 27, at 12:59 UTC, with a total exposure time of approximately 90 ks for each DU (see Table 1 and the light curve in Fig.
 1).
-Data have been processed for model-independent polarimetric analysis using the ixpeobssim package version 30.2.1 (Baldini et al., 2022).
-Spectral and spectro-polarimetric analysis was performed using HEASoft version 6.33 and the standard FTOOLS (Nasa High Energy Astrophysics Science Archive Research Center (2014), Heasarc) with the IXPE CALDB released on 28 February 2024.
+Data were processed for model-independent polarimetric analysis using the ixpeobssim package version 30.2.1 (Baldini et al., 2022).
+Spectral and spectropolarimetric analysis was performed using HEASoft version 6.33 and the standard FTOOLS (Nasa High Energy Astrophysics Science Archive Research Center (2014), Heasarc), with the IXPE Calibration Database (CALDB) that was released on 2024 February 28.
 Source photons were selected in a circular region with a radius of 100″ centered at the source position.
-As recommended by Di Marco et al.
-(2023b), given the brightness of the source, the background was not subtracted.
-The unweighted analysis is performed with ixpeobssim, while the weighted analysis is adopted for the spectropolarimetric analysis as suggested by Di Marco et al.
+As recommended by DiMarco_2023, given the brightness of the source, the background was not subtracted.
+The unweighted analysis was performed with ixpeobssim, while the weighted analysis was adopted for the spectropolarimetric analysis, as suggested by Di Marco et al.
 (2022a).
-Table 1: List of IXPE and Swift/XRT observations: Observation ID, start, stop, and the exposure times are reported.
-Observatory ObsID Start – Stop Exposure Time (s)
-Swift/XRT-XRT 00036688042 2024-02-26T05:05–2024-02-26T05:27 1366
+Table 1: List of IXPE and Swift/XRT observations.
+Observatory ObsID Start – Stop Instrument Exposure Time (s)
+Swift 00036688042 2024-02-26T05:05–2024-02-26T05:27 XRT 1366
 00036688043 2024-02-25T13:142024-02-25T17:59 892
@@
 00036688051 2024-02-27T11:05–2024-02-27T11:29 1443
-IXPE 03001101 2024-02-25T13:34–2024-02-27T12:59 90670
-90819
-90799
-Figure 1: The IXPE light curve obtained combining the 3 DUs (top panel) and the corresponding hardness ratio as a function of time (bottom panel).
+IXPE 03001101 2024-02-25T13:34–2024-02-27T12:59 DU1 90670
+DU2 90819
+DU3 90799
+Figure 1: IXPE light curve obtained by combining the three DUs (top panel) and the corresponding hardness ratio as a function of time (bottom panel).
 The data are binned in 200 s.
-The shaded regions mark the Swift/XRT observation periods.
-Dashed vertical lines illustrate the separation of the observation into five equal time bins of 9.5 h.
+Shaded regions indicate the Swift/XRT observation periods.
+Dashed vertical lines indicate the separation of the observation into five equal 9.5 h time bins.
 Figure 2: Hardness-intensity diagram obtained from IXPE data. The data are binned in 200 s.
 2.2 Swift/XRT
-The Neil Gehrels Swift Observatory (Gehrels et al., 2004) carrying three instruments on board, enables the most detailed observations of gamma-ray bursts to date; the X-ray Telescope (XRT) is one of the instruments on board based on a sensitive, flexible, autonomous X-ray CCD imaging spectrometer.
-Swift/XRT coordinated observations with IXPE for GX 13$+$1 were performed; they are used to monitor the status of the source and obtain spectral information during IXPE observation.
-Given the source brightness, the Swift/XRT observations were performed in Windowed Timing mode (WT).
+Carrying three instruments on board, the Neil Gehrels Swift Observatory (Gehrels et al., 2004) enables the most detailed observations of gamma-ray bursts to date; the X-ray Telescope (XRT), one of the instruments on board, is based on a sensitive, flexible, autonomous X-ray CCD imaging spectrometer.
+Swift/XRT-coordinated observations with IXPE for GX 13$+$1 were performed, and used to monitor the status of the source and obtain spectral information during IXPE observations.
+Given the source brightness, the Swift/XRT observations were performed in windowed timing (WT) mode.
 Ten Swift/XRT pointings covered the IXPE observation, as reported in Table 1 and Fig.
 1.
-Swift/XRT data were extracted using HEASoft v6.33 and standard FTOOLS Nasa High Energy Astrophysics Science Archive Research Center (2014) (Heasarc).
+Swift/XRT data were extracted using HEASoft v6.33 and standard ftools (Nasa High Energy Astrophysics Science Archive Research Center (2014), Heasarc).
 Source and background extractions were performed using Swift/XRT imaging capabilities in an annulus with an inner radius of about 5″ and an outer radius of about 60″ to avoid possible pile-up effects (Romano et al., 2006).
-The Swift/XRT data have been fitted in the energy band 0.7–8 keV to have sufficient statistics in the spectral fits and grouped to have at least 50 counts bin-1.
-In the analysis, the response matrices released in the HEASARC CALDB on 2023 July 25 have been applied.
+The Swift/XRT data were fitted in the energy band 0.7–8 keV to ensure sufficient statistics in the spectral fits and grouped to have at least 50 counts bin-1.
+Response matrices released in the HEASARC CALDB on 2023 July 25, were applied in this analysis.
 3 Data analysis
@@
 1 (top panel), the source rate was rather stable, with a slight increase at the end of the observation.
-The source hardness in the IXPE data appears to have no strong variations, remaining almost constant throughout the observation; this is in contrast to the previous IXPE observation, reported in B24, where a dip corresponding to a slight hardening was present.
-In Fig.
-2, the hardness intensity diagram obtained from the IXPE data is reported and shows that hardness did not change by more than 10%.
-Swift/XRT observations, covering with short snapshots the IXPE observation, allow us to monitor the source spectral state.
-GX 13$+$1 studies in the past showed a spectrum that can be well described, as typical in the atoll and Z sources, by a soft diskbb component describing disk emission and Comptonization emission from the spreading/boundary layer that can be described by a bbodyrad component (Díaz Trigo et al., 2012).
-Better modeling of the hard Comptonization component, e.g., applying compTT or compTB, is not possible due to the limited IXPE and Swift/XRT energy bands.
-In the literature, XMM–Newton spectra of GX 13$+$1 are reported showing wind and reflection features (Díaz Trigo et al., 2012), but neither IXPE nor Swift/XRT allow us to study these features, so in the following the simple spectral model tbabs*(diskbb+bbodyrad) will be applied.
+The source hardness in the IXPE data appears to have no strong variations, showing only a slight decrease with the count rate increase toward the end of the observation; this contrasts with the previous IXPE observation reported in B24, where a dip corresponding to a slight hardening was present.
+Figure 2, which reports the HID obtained from the IXPE data, shows that hardness did not change by more than 10%.
+Consisting of short snapshots during the IXPE observation, the Swift/XRT pointings enable us to monitor the spectral state of the source.
+Previous studies of GX 13$+$1 have shown that its spectrum can be well described, as is typical for atoll and Z sources, by a soft diskbb component representing disk emission and a bbodyrad component representing Comptonization emission from the spreading and/or boundary layer (Díaz Trigo et al., 2012).
+More detailed modeling of the hard Comptonization component, such as using compTT or compTB, is not possible due to the limited energy bands of IXPE and Swift/XRT.
+Literature on XMM–Newton spectra of GX 13$+$1 reports wind and reflection features (Díaz Trigo et al., 2012), but neither IXPE nor Swift/XRT allows us to study these features.
+Therefore, in the following analysis, we will apply the simplified spectral model, tbabs*(diskbb+bbodyrad).
 3.1 Spectroscopy
-Given the better spectral capabilities, we used Swift/XRT data as a driver for spectral modeling and to monitor spectral variations.
-The analysis is carried out in the energy band 0.7–8 keV, best-fit values are reported for each observation in Table 2 with uncertainties at the 68% confidence level (CL), and ObservationIDs are ordered by the observation date to show the time evolution of the spectral model.
-Table 2: Best-fit parameters for the spectral model TBabs*(diskbb+bbodyrad) as obtained by the different Swift/XRT spectra.
+Given its superior spectral capabilities, we used Swift/XRT data as the primary source for spectral modeling and to monitor spectral variations.
+The analysis was conducted in the 0.7–8 keV energy band, with best-fit values reported for each observation in Table 2, including uncertainties at the 68% confidence level (CL).
+Observation IDs are ordered by observation date to illustrate the time evolution of the spectral model.
+Table 2: Best-fit parameters for the TBabs*(diskbb+bbodyrad) spectral model as obtained by the different Swift/XRT spectra.
 […]
-$N_{\text{H}}$ (1022 cm2) $kT_{\text{in}}$ (keV) norm $kT_{\text{bb}}$ (keV) norm (10-9 erg cm-2 s-1)
+$N_{\text{H}}$ (1022 cm-2) $kT_{\text{in}}$ (keV) norm $kT_{\text{bb}}$ (keV) norm (10-9 erg cm-2 s-1)
 00036688043 $4.54^{+0.18}_{-0.14}$ $1.1^{+0.6}_{-0.2}$ $250^{+260}_{-160}$ $1.5^{+0.6}_{-0.1}$ $100^{+60}_{-90}$ 0.93 6.0 0.35
@@
 Errors correspond to the 68% CL.
-The reported flux has a typical uncertainty of the order of $0.3\times 10^{-9}$.
+The reported flux has a typical uncertainty on the order of $0.3\times 10^{-9}$.
 The hardness is defined as in Eq.
 (1).
-In the spectral analysis, the Swift/XRT data are fitted with the model: tbabs*(diskbb+bbodyrad).
-In the different pointings, it is possible to observe that the absorption is in average $\sim 4.7\times 10^{22}$ cm-2, while the disk temperature ranges between 0.7–1.1 keV, about the bbodyrad component – which in our case is used to describe a Comptonization emission from the boundary layer – the temperature ranges in the $\sim 1.3-1.8$ keV.
-The flux is in the range $\sim$(4–6)×$\times 10^{-9}$ erg cm-2 s-1 except in the last pointing where it reaches the highest value of $\sim 8\times 10^{-9}$ erg cm-2 s-1.
-The hardness, reported in Table 2 is defined as
-$\text{Hardness}=\frac{\text{Flux}_{0.7-3\leavevmode\nobreak\ \text{keV}}-\text{Flux}_{3-8\leavevmode\nobreak\ \text{keV}}}{\text{Flux}_{0.7-8\leavevmode\nobreak\ \text{keV}}}$ (1)
-and the spectral models report an almost constant hardness value except for the case of ObservationID 00036688045 where the flux is at a minimum and there is a slight hardening, but this Swift/XRT observation (third snapshot) is performed mainly during an IXPE occultation period.
-Thus, the IXPE data are not affected by this slight hardening, while the other snapshots confirm the IXPE result of an almost constant hardness ratio during observation.
-In the following analysis, for the spectro-polarimetric study, this ObservationID will be excluded, while the other ones will be used in the joint fit with the IXPE data.
-Taking into account the whole IXPE observation, and the Swift/XRT spectra, except the ObservationID 00036688045, we obtain a good joint spectrum model.
+In the spectral analysis, the Swift/XRT data were fitted using the tbabs*(diskbb+bbodyrad)model.
+Across the different pointings, the absorption column density was on average $\sim 4.7\times 10^{22}$ cm-2, while the disk temperature ranged between 0.7 and 1.1 keV.
+Regarding the bbodyrad component – which in our case is used to describe a Comptonization emission from the boundary layer – the temperature ranged from $\sim 1.3-1.8$ keV.
+The flux was typically in the range of $\sim$(4–6)×$\times 10^{-9}$ erg cm-2 s-1, except in the final pointing, where it reached the highest value of $\sim 8\times 10^{-9}$ erg cm-2 s-1.
+The hardness (see Table 2) defined as
+$\text{Hardness}=\frac{\text{Flux}_{3-8\leavevmode\nobreak\ \text{keV}}-\text{Flux}_{0.7-3\leavevmode\nobreak\ \text{keV}}}{\text{Flux}_{0.7-8\leavevmode\nobreak\ \text{keV}}}$ (1)
+has an almost constant value, except for Observation ID 00036688045, where the flux was minimal and the hardness maximal.
+However, this Swift/XRT observation (third snapshot) was conducted mainly during an IXPE occultation period.
+Thus, the IXPE data are not affected by this slight hardening, and the other snapshots confirm the IXPE result of an almost constant hardness ratio during the observation.
+In the following analysis for the spectropolarimetric study, this Observation ID is excluded, and the others are used in the joint fit with the IXPE data.
+Considering the entire IXPE observation and the Swift/XRT spectra, except the Observation ID 00036688045, we obtain a good joint spectrum model.
 The best-fit parameters are reported in Table 3 and Fig.
 3.
-To take into account the calibration uncertainties in the IXPE data, (see, e.g., Di Marco et al., 2022b), we left the gain slope and offset of the IXPE-DUs free to vary.
-Table 3: Best-fit parameters of the model const*tbabs*(diskbb+bbodyrad) for the joint Swift/XRT and IXPE spectra.
+To account for the calibration uncertainties in the IXPE data (see, e.g., Di Marco et al., 2022b), we allowed the gain slope and offset of the IXPE DUs to vary freely.
+Table 3: Best-fit parameters of the const*tbabs*(diskbb+bbodyrad) model for the joint Swift/XRT and IXPE spectra.
 Model Parameter Value
-tbabs $N_{\text{H}}$ (1022 cm2) $4.45_{-0.05}^{+1.0}$
+tbabs $N_{\text{H}}$ (1022 cm-2) $4.45_{-0.05}^{+1.00}$
 diskbb $kT_{\text{in}}$ (keV) $0.88_{-0.07}^{+0.10}$
@@
  [^2] Errors correspond to the 68% CL.
-Figure 3: Joint spectral fit Swift/XRT and IXPE in the 1–8 keV energy range. In the top panel the spectra of each telescope are reported in $EF_{E}$ units, in the bottom panel the residuals are shown.
-The joint fit confirms the result expected for the average model from several Swift/XRT snapshots.
-In the following for the spectro-polarimetric analysis, the spectral model will be frozen to the present one.
-The IXPE gain correction is at level of 0.95 for the slope in all the three DUs and the offset in the range 70–100 eV.
+Figure 3: Joint spectral fit of Swift/XRT and IXPE data in the 1–8 keV energy range.
+In the top panel, the spectra of each telescope are reported in $EF_{E}$ units; residuals are shown in the bottom panel.
+The joint fit confirms the expected result for the average model from several Swift/XRT snapshots.
+In the following spectropolarimetric analysis, the spectral model is fixed to the present one.
+The IXPE gain correction is at a level of 0.95 for the slope in all three DUs, with the offset in the range of 70–100 eV.
 3.2 Polarimetry
-Spectroscopy revealed that neither the light curve nor the hardness have strong dips or features that would suggest separating the observation into several parts to be studied independently.
+The spectroscopy reveals that neither the light curve nor the hardness has strong dips or features that would suggest the observation should be separated into several parts to be studied independently.
 To study the time variability of the polarimetric properties, we followed the same approach as in B24 and split this new IXPE observation into five equal time bins spanning 9.5 h.
@@
 Figure 4: Time dependence of the normalized Stokes parameters $q$ and $u$ (top), and of the PD and PA (bottom) obtained by the pcube algorithm using the data separated into 9.5 h time bins.
-The confidence regions are reported at $68\%$.
-Figure 5: Energy resolved normalized Stokes parameters $q$ and $u$ (top), and PD and PA (bottom) obtained by the pcube algorithm using the reported energy bins for the time-averaged data.
+Confidence regions are reported at $68\%$.
+Figure 5: Energy-resolved normalized Stokes parameters $q$ and $u$ (top), and PD and PA (bottom) obtained by the pcube algorithm using the reported energy bins for time-averaged data.
 The confidence regions are reported at $68\%$ CL.
-The vectors correspond to the best-fit polarization parameters of the diskbb and bbodyrad components, respectively, as obtained in the spectro-polarimetric analysis and reported in Table 4.
+The vectors correspond to the best-fit polarization parameters of the diskbb and bbodyrad components, respectively, as obtained using spectropolarimetric analysis and reported in Table 4.
 […]
-3.3 Spectro-polarimetry
+3.3 Spectropolarimetry
 […]
-Table 4: IXPE spectro-polarimetric fit of the Stokes parameters $I$, $Q$, and $U$ using the spectral model from Table 3.
-Model PD/$A_{1}$ (%) $A_{\rm slope}$ (keV-1) PA/$\psi_{1}$ (deg) $\psi_{\rm slope}$ (deg,keV-1) $\chi^{2}$/d.o.f.
+Table 4: IXPE spectropolarimetric fit of the Stokes parameters $I$, $Q$, and $U$ using the spectral model from Table 3.
+Model PD/$A_{1}$ (%) $A_{\rm slope}$ (keV-1) PA/$\psi_{1}$ (deg) $\psi_{\rm slope}$ (deg keV-1) $\chi^{2}$/d.o.f.
 tbabs*(diskbb+bbodyrad)*polconst $2.4\pm 0.3$ $28\pm 3$ 243/250=0.97
@@
 The left, middle, and right panels are for the Stokes parameters $I$, $Q$, and $U$, respectively.
-The fit is performed in the 2–8 keV energy band using the three IXPE detectors and applying the model tbabs*(diskbb*polconst+bbodyrad*polconst).
-The total model is shown with the solid black line, while the diskbb and bbodyrad with the dashed and dotted lines, respectively.
-The lower sub-panels show the residuals between the data and the best fit.
-In the joint spectral fit, we apply a rebinning to have at least 50 counts per bin as in Swift/XRT, while for the spectro-polarimetric analysis, a constant grouping is applied to have energy bins of 200 eV.
+The fit is performed in the 2–8 keV energy band using the three IXPE detectors and applying the tbabs*(diskbb*polconst+bbodyrad*polconst) model.
+The total model is shown with the solid black line, while diskbb and bbodyrad are indicated with the dashed and dotted lines, respectively.
+The lower subpanels show the residuals between the data and the best fit.
+In the joint spectral fit, we applied a rebinning to obtain at least 50 counts per bin as in Swift/XRT, while for the spectropolarimetric analysis, a constant grouping was applied to have energy bins of 200 eV.
 4 Discussion
-Although interesting as a single observation, the results reported in this article provide the most information when compared to the results of the October observations reported in B24.
-Overall, in October IXPE measured a PD of $\sim$1.4and a PA of $\sim-2\degr$, significantly different from the current results.
+Although intriguing on its own, the observation reported in this article yields the most insight when compared to the results of the October 2023 observations reported in B24, when we measured an overall PD of $\sim$1.4and a PA of $\sim-2\degr$, significantly different from the current results.
 However, the reason for the depolarization in the first observation was fast variability in the PA.
@@
 7 and compared to the results of the current observation.
-We note that the current observation is similar in PA to the post-dip state of B24.
-The lower PD in the entire current observation most probably comes from averaging the entire observation in time and energy, as for the previous one.
-In Fig.
-4, the Bin 3, for instance, shows a polarization within errors similar to the one measured after the dip in B24.
+We note that the current observation is similar in PA to the post-dip state of B24, but the PD is lower probably due to the averaging over time and energy.
+Bin 3 in Fig.
+4, for instance, shows a polarization within errors similar to the one measured after the dip in B24.
 […]
-Both HID and the shape of the spectrum confirm that GX 13$+$1 was in a soft state throughout the observation.
+Both the HID and the shape of the spectrum confirm that GX 13$+$1 was in a soft state throughout the observation.
 We do not see any significant change of state in the IXPE and Swift/XRT energy bands, although, in the last part of the observation, the flux rises, which is a sign of the source going into an even softer state.
-If we look at the HID of the source in the October observation (B24, Figure 2), we see that the source is almost in the same state, except for the dipping period.
-In this new observation, since simultaneous Swift/XRT observations offer a better spectral capability than IXPE, the spectral modeling and decomposition into two main components are better determined, allowing for a better spectro-polarimetric analysis.
+If we look at the HID of the source in the October observation (Fig.
+2 in B24), we see that the source is almost in the same state, except for the dipping period.
+During the new observation period, the spectral modeling and decomposition into two main components are better determined, since simultaneous Swift/XRT observations offer a better spectral capability than IXPE, thereby allowing for a better spectropolarimetric analysis.
 The light curve does not show any peculiar variability: unlike the one from the October observation, the current one has no strong dips or rapid changes, and neither does the hardness ratio.
-Despite this constant hardness, we attempted to study the possible variation in 9.5 h equal time bins similarly to B24.
-This analysis reports a hint for a variation in PD in Fig.
-4, in particular is interesting that in Bin 5, corresponding to a softening of GX 13$+$1 there is a lower polarization and an indication for a possible rotation with respect to the other 4 time bins.
-The clear rotation of PA observed during the first observation is not present this time, but the PA appears to have remained aligned with the post-dip polarization measured in B24.
-If we compare the positions of the points on the $(q,u)$ plane with those obtained from the October 2023 observation (Figure 7 in B24), we see a different pattern, but the points are still in the I or IV quadrant.
-The biggest difference is in order: in October, we saw the points following the straight line path, but here the variability shows a different pattern.
+Despite this constant hardness, we attempted to study the possible variation in 9.5 h equal time bins similar to B24.
+This analysis suggests possible variation in PD in Fig.
+4; in particular, it is interesting to note that in Bin 5, corresponding to a softening of GX 13$+$1, there is lower polarization and an indication of a possible rotation with respect to the other four time bins.
+The clear rotation of the PA observed during the first observation is not present this time, but the PA appears to have remained aligned with the post-dip polarization measured in B24.
+If we compare the positions of the points on the $(q,u)$ plane with those obtained from the October 2023 observation (Fig.
+7 in B24), we see a different pattern; however, the points are still in the I or IV quadrant.
+The most significant difference lies in the pattern: in October, we observed the points following a straight line path, whereas here, the variability exhibits a different pattern.
 […]
 From the aligned systems, we expect the polarization vectors of the emission coming from the disk and the Comptonized component to be either nearly parallel or perpendicular to each other.
-Hence, we expect some misalignment in the system that could support this difference in the PA.
-Another thing to notice is a very high PD of the disk component: as was mentioned in Loktev et al.
-(2022), we do not expect a polarization higher than 4% in the IXPE energy range even for the highest inclination.
-This value is compatible with the spectro-polarimetric results at 90% CL.
-However, knowing that the source is likely to have an inclination lower than 80°(Díaz Trigo et al., 2012; Tomaru et al., 2020), we need to add the scattering of the disk emission in the wind to reach the PD of 6.5%.
-We assume that the Comptonized component comes from the boundary layer, as it is expected to align in PA with the disk emission, and the difference in PAs is less than 45°.
-In search of a source of misalignment, we looked at the previous observations of IXPE.
-It is possible (although unexpected) that, similar to the Cir X-1 (Rankin et al., 2024), the rotation axis of the neutron star is slightly misaligned from the rotation axis of the binary system.
-The boundary layer would then be dragged out of the disk plane by the motion of the star.
-Another option would be the peculiar geometry of the wind in the system.
-Finally, we want to address the significant difference between the two observations of GX 13$+$1.
-The object was in the same spectral state but half an orbit away from the previous observation, when the source clearly showed a rotation of the PA with time.
-However, in February 2024, there was no clear monotonic change in the PA, while we still see variations in the PA.
-One possibility is that the mechanism responsible for the dip in the light curve during the first observation was also responsible for the rotation of the PA.
-However, this mechanism is unknown.
-What we can address is that from this new observation, it appears that the PA remained aligned with the one observed in the end of the previous observation.
-Then, at the end of the present IXPE pointing, when the flux grows and the source goes into a softer state, the PD decreases and rotates to a PA similar to the one we observed in October before the dip.
-This could suggest, similarly to what we observed in Cir X-1, a variation of the PA with the state of the source.
-This could mean a variation of the geometry of the hot region in different states; this can also be confirmed by the indication we obtained in the October 2023 observation when the energy dependence of the polarization changed before and after the dip going from an energy-dependent behavior, as observed e.g.
-for 4U 1820--30 (Di Marco et al., 2023a), into a constant polarization with energy, as observed, e.g., for Sco X-1 in La Monaca et al.
+Hence, we expect some misalignment in the system that would account for this difference in the PA.
+Additionally, it is important to note the exceptionally high PD of the disk component.
+As stated by Loktev et al.
+(2022), polarization exceeding 4% in the IXPE energy range is not anticipated, even at the highest inclination.
+This value is compatible with the spectropolarimetric results at 90% CL.
+However, knowing that the source is likely to have an inclination lower than 80°(Díaz Trigo et al., 2012; Tomaru et al., 2020), we need to add the scattering of the disk emission in the wind to reach a PD of 6.5%.
+We assumed that the Comptonized component comes from the boundary layer, as it is expected to align in PA with the disk emission.
+The observed difference in PAs is less than 45°.
+In search of a source of misalignment, we examined previous observations by IXPE.
+It is possible, though unexpected, that the rotation axis of the neutron star is slightly misaligned from the rotation axis of the binary system, similar to the Cir X-1 (Rankin et al., 2024).
+In such a scenario, the boundary layer could be dragged out of the disk plane by the motion of the star.
+Another potential explanation is the peculiar geometry of the wind within the system.
+Finally, we wish to address the significant difference between the two observations of GX 13$+$1.
+Although the object was in the same spectral state, it was observed half an orbit away from the previous observation, during which the source clearly exhibited a rotation of the PA over time.
+However, in February 2024, there was no evident monotonic change in the PA, despite ongoing variations.
+One possibility is that the mechanism responsible for the dip in the light curve during the first observation also caused the rotation of the PA.
+However, this mechanism remains unidentified.
+Notably, the recent observation suggests that the PA remained aligned with that observed at the end of the previous observation.
+As the flux increased and the source transitioned to a softer state at the end of the present IXPE pointing, the PD decreased and rotated to a PA similar to that observed in October before the dip.
+This pattern might indicate, as observed in Cir X-1, a variation in the PA corresponding to the state of the source.
+This implies a potential variation in the geometry of the hot region across different states.
+This hypothesis is supported by the October 2023 observation, where the energy dependence of the polarization shifted from an energy-dependent behavior, as seen in 4U 1820--30 (Di Marco et al., 2023), to a constant polarization with energy, as observed, for example, in Sco X-1 in La Monaca et al.
 (2024).
-Present data do not allow for a clear statement on this, but we hope in the future we could obtain further observations to improve the significance for a detailed analysis of the polarization along the states of atoll and Z sources.
-A further point is the contribution of the wind, as we are unaware of the wind behavior during both observations, we can also assume that it could be responsible for the differences in polarimetry, yet we have no way to support this claim.
+While current data do not allow for a definitive conclusion, we hope that future observations will provide sufficient significance for a detailed analysis of the polarization across different states of atoll and Z sources.
+Additionally, the role of the wind remains an open question.
+Given our lack of knowledge regarding wind behavior during both observations, it is plausible that the wind could contribute to the observed differences in polarimetric properties, although we currently have no means to substantiate this hypothesis.
 5 Summary
@@
 In this article, we report the highly significant detection of polarization from this source in the soft state.
-Spectroscopic analysis showed that the spectrum is close to that in Díaz Trigo et al.
-(2012), and the hardness of the emission confirmed the lower left banana state, as well as the further softening of the spectrum at the end of this new observation.
+Spectroscopic analysis shows that the spectrum is close to that in Díaz Trigo et al.
+(2012), and the hardness of the emission confirms the lower left banana state, as well as the further softening of the spectrum at the end of this new observation.
 We used a diskbb model for the softer disk emission, and the bbodyrad model to approximate the harder component associated with the spreading or boundary layer.
-The polarimetric analysis provided the overall PD of the source during the observation at 2.5% at ¿$5\sigma$ CL, and a PA at 24°.
-We also saw a marginally significant decrease in PD and PA with energy.
-The same trend was confirmed in the spectro-polarimetric analysis, which showed that the softer component is stronger polarized than the harder component with the PA differing by $\sim$40°.
-The results of the spectro-polarimetric analysis using xspec performed under the assumption of one overall polarization model agreed with the pcube polarimetric results.
+The polarimetric analysis provided the overall PD of 2.5% at ¿$5\sigma$ CL and a PA of 24°.
+We also see a marginally significant decrease in PD and PA with energy.
+The same trend was confirmed by the spectropolarimetric analysis, which showed that the softer component was more strongly polarized than the harder component, with the PA differing by $\sim$40°.
+The results of the spectropolarimetric analysis using xspec performed under the assumption of one overall polarization model were in agreement with the pcube polarimetric results.
 We studied the differences and similarities between the first (B24) and the second observations of GX 13$+$1 by IXPE.
-Our main conclusion is that we saw a significantly higher overall polarization in the second observation, even though the source was in the same state during both observations.
-The depolarization comes from the variability of polarimetric properties, which was larger in the first observation.
-We discussed possible reasons for the different polarimetric behavior in the two observations.
+Our main conclusion is that the overall polarization is significantly higher in the second observation despite the source being in the same state during the two observations.
+Depolarization was caused by the variability in polarimetric properties, which was larger during the first observation.
+We have discussed possible reasons for the different polarimetric behavior in the two observations.
 Acknowledgements.
@@
 This research has been supported by the Finnish Cultural Foundation grant 00240328 (AB) and the Academy of Finland grant 333112 (AB, JP, SVF, VL).
-ADM, FLM are partially supported by MAECI with grant CN24GR08 “GRBAXP: Guangxi-Rome Bilateral Agreement for X-ray Polarimetry in Astrophysics”.
+ADM and FLM are partially supported by MAECI with grant CN24GR08 “GRBAXP: Guangxi-Rome Bilateral Agreement for X-ray Polarimetry in Astrophysics”.